OpenStudy (anonymous):

x + (3x + 2) +1/3 x = 29........help please?

OpenStudy (anonymous):

118 1/3 is your answer

OpenStudy (australopithecus):

x + (3x + 2) +1/3 x = 29 first off you dont need the brackets because you are not multiplying anything into them \[ x + 3x + 2 + \frac{1}{3} x = 29\] can you solve it now?

OpenStudy (australopithecus):

or at least simplify it?

OpenStudy (australopithecus):

a bit more

OpenStudy (anonymous):

no lol sorry im confused

OpenStudy (australopithecus):

Well the next step is to collect like terms, and isolate x on one side of the equation \[ 4x + 2 + \frac{1}{3}x - 2 = 29 - 2\]

OpenStudy (australopithecus):

do you follow?

OpenStudy (anonymous):

omg that is like a bunge of confusion to me...

OpenStudy (australopithecus):

Remember the rule that as long as you do the same thing to both sides of an equation you maintain the equality for example, 2 = 2 I can subtract -2 from both sides and it remains equal 2 - 2 = 2 - 2 0 = 0 so by using this rule I'm isolating the x values on one side of the equation also, remember x = a number, it can be any number for example 1, 2, 3 thus if you have 3x you have, x + x + x so you have 3*x if you added another x to 3x than you would have x + x + x +x = 4x you would have 4x or 4*a number

OpenStudy (australopithecus):

isolating the x value on one side of the equation is essential to determining what x is equal to

OpenStudy (australopithecus):

do you follow still?

OpenStudy (anonymous):

itll take a secound lol maybe

OpenStudy (australopithecus):

no problem take your time, it is important you understand this. If there is anything you need explained more clearly just ask and I will try to improve my explanation.

OpenStudy (anonymous):

okie dokie

OpenStudy (australopithecus):

Correction * remember x = a number, it could be any number for example 1, 2, or 3 So just say x = 3 and you have the equation 3x + x we can sub in 3 to solve it because we know x =3 3(3) + 3 = 3 + 3 + 3 + 3 = 9 + 3 = 12 as you can see x is just a number

OpenStudy (australopithecus):

tell me when you understand the manipulations I preformed

OpenStudy (anonymous):

okay(:

OpenStudy (australopithecus):

give up?